OpenStudy (anonymous):

What is wrong with this Domain and Range? y = -2x -5 {xeI / x<-15/2 , x>-2} {yeI / y>10 , y<-1}

OpenStudy (anonymous):

Er.. the domain is \(x\in \mathbb{R}\) as is the range?

OpenStudy (anonymous):

the range was the restriction that the teacher gave us and it started with \[y \epsilon \]

OpenStudy (anonymous):

I mean y is an element of I

OpenStudy (anonymous):

All right, using that, then, for some domain: \[ x\in(a,b) \]We must have a range of a linear function of: \[ f(x)\in\left(\;\min(f(a), f(b)),\; \max(f(a), f(b)\right)\;) \]

OpenStudy (anonymous):

wait what does the I even mean?

OpenStudy (anonymous):

I don't know... l? That's not stated in the problem...?

OpenStudy (anonymous):

I'm not sure if it means y is an element of integers or something like that... but my teacher said my calculations were right except that I got something fundamentally wrong with my domain...

OpenStudy (anonymous):

No, the integers is \(\mathbb{Z}\), and your region is correct. Just that, you cannot say that \(x,y\in \mathbb{I}\), because \(\mathbb{I}\) typically stands for an identity function, not the integers. Instead, you should say: \(x,y\in \mathbb{R}\), which means reals.

OpenStudy (anonymous):

so for my domain it should be x∈R ?

OpenStudy (anonymous):

OH. I think I get what your question is. So you were given the constrains for \(y\) to be an integer and had to find the fitting \(x\)? Am I incorrect? In such a case, yes there is something fundamentally wrong in your answer, as \(-\frac{15}{2}\) is not an integer.

OpenStudy (anonymous):

yes the range was the constraint and I had to find the domain. So does y∈I mean that I need to work with integers only?

OpenStudy (anonymous):

so do I have to round it?

OpenStudy (anonymous):

\(y\in \mathbb{I}\) does not exist. \(\mathbb{I}\) is not a real field.

OpenStudy (anonymous):

soo what does that mean for the domain LOL

OpenStudy (anonymous):

there is no solution?